OpenStudy (anonymous):

Prove recurrence relation solution by induction: Initial recurrence relation: \(a_n = a_{n-1} + 1 + 2^{n-1}; a_0 = 0\) Solution: \(a_n = n + 2^n - 1\)

OpenStudy (anonymous):

oh it was right?

OpenStudy (anonymous):

Seemingly so! We won't know until we prove it though ;)

OpenStudy (anonymous):

assume \[a_k=k+2^k-1\] prove \[a_{k+1}=(k+1)+2^{k+1}-1\] aka \[a_{k+1}=k+2^{k+1}\]

OpenStudy (anonymous):

Good start...

OpenStudy (anonymous):

it is mostly algebra since \[a_{k+1}=a_k+1+2^{k-1}\] and you can replace \(a_k\) by \(k+2^k-1\) by induction

OpenStudy (anonymous):

you get it right away

OpenStudy (anonymous):

i don't get it :(

OpenStudy (anonymous):

do you get that by induction \[a_{k+1}=k+2^k-1+1+2^{k-1}\]

OpenStudy (anonymous):

no i don't know induction

OpenStudy (anonymous):

\[a_{k+1}=\overbrace{k+2^k-1}^{a_k}+1+2^{k-1}\]

OpenStudy (anonymous):

we replace \(a_k\) by the formula we assume to be true

OpenStudy (anonymous):

ok then what

OpenStudy (anonymous):

algebra to show that \[a_{k+1}=(k+1)+2^{k+1}-1\] which is the formula with \(n\) replaced by \(k+1\)

OpenStudy (anonymous):

it is not so easy to explain the idea behind induction here in a chat box, i am tempted to say that if you do not know how to do a proof by induction then you cannot do it but what i wrote is what you have to do, and it comes down to algebra in showing that \[a_{k+1}=\overbrace{k+2^k-1}^{a_k}+1+2^{k-1}=(k+1)+2^{k+1}-1\]

OpenStudy (anonymous):

ok then where or how can it be explained to me? you know of any reliable resources on the web that might drive the point home?

OpenStudy (anonymous):

it is probably in your text, or you can google it first you show it is true for \(n=1\) then you assume it is true for \(n=k\) and using that assumption prove it is true for \(n=k+1\) there are piles of youtube vidoes usually dealing with summation formulas

OpenStudy (anonymous):

alright i'll try to figure it out thank you

OpenStudy (anonymous):

yw